# Generalized Euler Equation from Effective Action: Implications for the Smarr Formula in AdS Black Holes

### v1

@@ 3.2 Generalized Euler equation
To generalize the Euler equation, the only viable approach is to introduce a new variable into the effective action for perfect fluids. This new variable must be a diffeomorphism-invariant scalar with respect to $g_{ab}$ and gauge-invariant under the symmetry transformations defined in (21) and (22). As discussed at the end of section 2, there are situations where the standard scaling argument for deriving the Euler equation breaks down—such as when the system’s volume is compact. In this case, the radius of the compact space introduces an additional geometrical scale to the physical problem. This suggests that the new variable in the effective action should have a geometrical origin. The gauge-invariant quantity that satisfies these requirements is
@@
The definition of this variable closely resembles the definition of the chemical potential $\mu$ presented in (31). However, the temporal scalar field $\sigma^{0}$, which represents the clock for the fluid element, differs in nature from the Stueckelberg field $\varphi$ associated with U(1) symmetry [^3] The variable $y$ has been previously studied in the literature, but Ballesteros did not clearly distinguish between the temporal scalar field $\sigma^{0}$ and the gauge scalar field $\varphi$.. Given the definition of $y$, its significance may not be immediately clear. Therefore, let us first derive the generalized Euler equation and subsequently discuss the physical meaning of $y$.
@@
Inside the parenthesis of the last term, we recognize the local pressure $p$ given in (42). Therefore, we obtain a new relation among thermodynamic variables:
@@
This expression is the generalized Euler equation that we were seeking. The additional correction involves a derivative of the pressure $p$ with respect to $y$. Therefore, if the thermodynamics of a fluid in a given spacetime does not respond to variations in $y$, the extra correction $\frac{\partial p}{\partial y}$ vanishes. Another equivalent way of writing (47), which may be useful for clarity, is as follows:
@@
where $\omega$ is the density of the grand canonical potential. This form clearly illustrates that the additional scale in the problem, parameterized by $y$, contradicts the conventional wisdom that, in the thermodynamic limit, the grand canonical potential $\omega$ is proportional to the pressure $p$.
To understand the meaning of $y$, let us consider a scenario of global thermal equilibrium, where it is possible to align the fluid clock with physical time, i.e., $\sigma^{0}=x^{0}$. In this case, we obtain $y=u^{0}$, an expression that may be challenging to interpret. To gain some intuition, let us perform a Weyl transformation of the physical metric $g^{\prime}=Wg$, with $W$ being a conformal factor. Under this transformation, the fluid velocity changes as $u^{\prime a}=\frac{u^{a}}{W}$. Consequently, we arrive at the following relation:
$y=u^{\prime 0}=\frac{u^{0}}{W}~{}.$ (49)
This relation suggests that the variable $y$ is associated with the inverse of the conformal factor $W$. Instead of performing an explicit Weyl transformation, we can express the physical metric $g$ as a conformal factor times a reference metric $\bar{g}$, i.e., $g=W\bar{g}$.
It is important to emphasize that the motivation for introducing the variable $y$ was to capture an additional scale that a given spacetime could introduce into the thermodynamics of a fluid. This is why, we should not expect a microscopic interpretation for the variable $y$, unlike the other variables in the new thermodynamic potential $\chi=\chi(s,\mu,y)$.
@@ 4 Smarr formula as a particular case of the generalized Euler equation
In this section, we extend the discussion from section 2 by examining the AdSd Reissner-Nordström black hole with arbitrary horizon topology. We adopt Eddington-Finkelstein coordinates, as this gauge simplifies the analysis of Weyl transformations compared to the Fefferman-Graham gauge. Moreover, we restrict the boundary dimension $(d-1)$ to be odd; otherwise, the conformal anomaly would need to be accounted for in the effective action. The line element is:
@@
$M+pV=TS+\mu Q+\frac{kA_{d-2}}{8\pi G_{N}}r^{d-3}_{h}~{}.$ (54)
For for $k=0$ the last term vanishes, thereby recovering the standard Euler equation. This is because, in flat space, there are no additional scales to consider.
We will now compute the thermodynamic potential $\chi$ (40) which is associated with the effective action for the holographic perfect fluid dual to the AdSd Reissner-Nordström black hole. Our objective is to place the holographic fluid in a boundary spacetime that exhibits spacetime dependence, allowing us to mimic a state that is near out of equilibrium. As a consequence, the thermodynamic variables now will become local local functions of the spacetime. To obtain the boundary metric with the desired properties, we perform the following radial diffeomorphism in the bulk line element (50):
@@
This line element describes a non-homogeneous, time-dependent background in which the fluid resides. The radial diffeomorphism (55) consequently induces a Weyl transformation on the boundary metric, leading to the following expression for the fluid velocity:
$u_{a}=\frac{W}{l}\delta^{a}_{0}~{}.$ (59)
@@
$ds^{2}_{fluid}=-0\cdot du^{2}+r^{2}_{+}d\Sigma^{2}_{k}~{},$ (60)
that matches the horizon metric (with no conformal factor). Given the target metric and the fluid velocity, we can use the formulas (29), (31), and (39) to compute the entropy density $s$, the chemical potential $\mu$, and the variable $y$, respectively:
$s=\frac{\Sigma_{k,d-2}}{4G_{N}}\frac{r_{h}^{d-2}}{W^{d-2}}~{},\quad\mu=\frac{\tilde{q}}{\gamma lr_{h}^{d-3}}\frac{1}{W}~{},\quad y=\frac{l}{W}~{},$ (61)
which are the variables that the thermodynamic potential $\chi$ depends on. We will utilize these variables to derive the explicit expression for the new thermodynamic potential $\chi$ from the energy density $\epsilon$ using equation (42). Our next step involves computing the boundary stress tensor via the formula (12) and the boundary metric given in (58). The components of the stress tensor are
$T^{t}_{\;t}=\frac{M}{W^{d-1}\Sigma_{k,d-2}}~{},$ (62)
$T^{\phi_{j}}_{\;\phi_{j}}=\frac{M}{(d-2)W^{d-1}\Sigma_{k,d-2}}~{}.$ (63)
@@
$\epsilon=\frac{(d-2)\Sigma_{k,d-2}}{16\pi G_{N}l^{2}}\left[\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l^{2})^{2}\mu^{2})+\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-1}{d-2}}\right]~{},$ (64)
where $M$ has been expressed in terms of $r_{h}$ as an intermediate step, and then in terms of the variables $s$, $\mu$, and $y$ given in equation (61). One can verify that the conformal equation of state: $p=\frac{\epsilon}{d-2}$ is satisfied.
@@
$\chi=\frac{(d-2)\Sigma_{k,d-2}}{16\pi G_{N}l^{2}}\left[-\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l^{2})^{2}\mu^{2})+\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-1}{d-2}}\right]~{}.$ (66)
@@
$T=\frac{(d-1)}{4\pi l^{2}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{1}{d-2}}+\frac{(d-3)(ky^{2}-(\gamma l^{2})^{2}\mu^{2})}{4\pi l^{2}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{-1}{d-2}}~{},$ (67)
$q=\frac{(d-2)\Sigma_{k,d-2}\gamma^{2}l^{2}\mu}{8\pi G_{N}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}~{}.$ (68)
@@
$\frac{\partial p}{\partial y}=k\frac{\Sigma_{k,d-2}y}{8\pi G_{N}l^{2}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}~{}.$ (69)
Thus, this term vanishes for a charged black brane, which is an expected result due to the absence of any extra geometrical scale arising from flat spacetime. By performing some straightforward algebra, we can verify that the generalized Euler equation holds:
@@
In conclusion, we have demonstrated that the Smarr formula is a particular instance of the generalized Euler equation derived in section 3.2. This finding suggests that holographic fluids adhering to the Smarr formula are not fundamentally different from other more conventional fluids. The key takeaway is that when an extra scale emerges from the geometry in which the fluid or thermal system resides, the standard scaling argument used to derive the Euler equation no longer applies, necessitating a generalization of this equation. This generalization was provided in section 3.2, where we derived the generalized Euler equation (70). In cases with a compact volume, the correction term $y\frac{\partial p}{\partial y}$ in the Euler equation can be interpreted as a finite-size effect.
5 Comments on extended black hole thermodynamics
In the previous sections, we argued and demonstrated that the Smarr formula is a specific case of the generalized Euler equation, particularly in the context of black holes in AdS. In this sense, the Smarr formula naturally fits within the framework of the AdS/CFT correspondence, with its origins rooted in well-established physics. This conclusion is supported by the fact that we derived the generalized Euler equation using effective field theory techniques applied to fluids. Only after establishing this result did we invoke the AdS/CFT correspondence to interpret the Smarr formula.
@@ 5 Comments on extended black hole thermodynamics
where $d$ is the dimension of the bulk spacetime [^4] For reviews about extended black hole thermodynamics see Mann4 ; Mann5 . Other relevant literature in the subject Mann2 ; Mann3 ; Clifford2 .. Our understanding of the results in Karch suggests that this identification is problematic within the context of holography.
@@
The main result of this paper—that the Smarr formula is a specific instance of the generalized Euler equation—contradicts this new proposal for extended black hole thermodynamics. The reasoning is as follows: perfect fluids, a particular case of hydrodynamics, differs fundamentally from renormalization group (RG) physics. While the number of degrees of freedom $N$ can indeed change with the energy scale in RG flow (as is well understood in the context of AdS/CFT, see Ioannis1 ; Ioannis2 ), this cannot explain the Smarr formula. In a perfect fluids, thermodynamic variables must be functions of spacetime to describe local thermal equilibrium, something that the number of degrees of freedom $N$ cannot do. The point is that we should not confuse a flow in a physical space where the fluid is on with a flow in theory space where an RG flow takes place. Finally, it is crucial to recognize that global equilibrium is merely a special case of local equilibrium. As a result, there is no justification for treating $C\sim N^{2}$ as a thermodynamic variable.
@@ 6 Conclusions and future directions
In this work, we derived a generalized Euler equation (47) using effective field theory techniques, granting our result a broader range of applicability independent of the AdS/CFT correspondence. We then applied the generalized Euler equation to a specific case: a holographic fluid derived from the boundary stress tensor associated with an AdSd Reissner-Nordström black hole, from which we recovered the Smarr formula. Based on these findings, we further argued that the framework of extended black hole thermodynamics cannot be physically correct. This is because any new thermodynamic variable that one might attempt to associate with the AdS radius $l$ cannot become a function of the boundary spacetime, as required for local equilibrium.
@@
It has been demonstrated that evaluating the on-shell action of spacetimes with a future event horizon (featuring a planar horizon topology) can reproduce the standard effective action presented in (32). This was achieved by setting up a double-Dirichlet boundary condition to evaluate the on-shell action Natalia1 ; Liu1 . It would be intriguing to extend this construction to the more general effective action (40) discussed in our article. The primary challenge lies in the fact that the new thermodynamic potential $\chi=\chi(s,\mu,y)$ does not coincide with the grand canonical potential $\omega=\omega(T,\mu,y)$, as might typically be expected from the evaluation of the Euclidean on-shell action.
@@
Another intriguing avenue for exploration would be the explicit construction of the thermodynamic potential $\chi$ for rotating black holes, which also exhibit a boundary stress tensor with a perfect fluid form Clifford0 ; Clifford . In Section 4, we focused on odd bulk dimensions $d$ to sidestep the complications of conformal anomalies that arise when the fluid resides in curved spacetimes with even boundary dimensions $(d-1)$. It would be valuable to develop the most general effective action for perfect fluids that can incorporate conformal anomalies. An example of this is the Schwarzian action, which can be interpreted as an effective action for an anomalous perfect fluid in 1D Jensen2 . A similar result applies to the boundary action of a non-rotating BTZ black hole with boundary sources turned on Jensen3 ; Kevin .
The Smarr formula remains valid in gravity theories with higher-order curvature corrections Tomas1 . This raises an important question: can the generalized Euler equation (47) accommodate these theories? If not, it would suggest that the Smarr formula once again points to the need for further generalizations of the Euler equation. In such a scenario, it is unclear to the author what the nature of the new variables introduced in the effective action might be. This uncertainty arises because the variable $y$ in (39) had a clear geometric motivation, making it challenging to envision new variables with equally well-defined physical interpretations.
@@
The isometries of the target metric $G_{MN}$, as given in (24), correspond to the emergent low-energy gauge symmetries of the effective action for hydrodynamics. More broadly, horizon asymptotic symmetries ASG1 ; ASG2 ; ASG3 ; ASG4 ; ASG5 can be interpreted as these emergent low-energy gauge symmetries in the boundary theory. Therefore, any generalization of horizon symmetries implies the existence of new gauge symmetries for holographic effective actions. Investigating the connection between this work and the new symmetries identified in Liu5 would be an intriguing direction for future research.

### v2

@@ 3.2 Generalized Euler equation
To generalize the Euler equation, we introduce a new variable into the effective action for perfect fluids. This variable must be a diffeomorphism-invariant scalar with respect to $g_{ab}$ and must remain gauge-invariant under the symmetry transformations defined in (21), (22). Furthermore, as discussed at the end of Section 2, there are scenarios in which the standard scaling argument to derive the Euler equation fails, particularly when the volume of the system is compact. In such cases, the compact space’s radius introduces an additional geometric scale to the problem. This observation suggests that the new variable in the effective action should have a geometric origin.
The gauge-invariant quantity that satisfies these criteria is
@@
The definition of this variable closely resembles the definition of the chemical potential $\mu$ presented in (31). However, the temporal scalar field $\sigma^{0}$, which represents the clock for the fluid element, differs in nature from the Stueckelberg field $\varphi$ associated with U(1) symmetry [^3] The variable $y$ has been previously studied in the literature, but Ballesteros did not clearly distinguish between the temporal scalar field $\sigma^{0}$ and the gauge scalar field $\varphi$.. Given the definition of $y$, its physical meaning may not be immediately clear. Therefore, let us first derive the generalized Euler equation and subsequently discuss the physical meaning of $y$.
@@
In parentheses of the last term, we recognize the local pressure $p$ given in (42). Therefore, we obtain a new relation among thermodynamic variables:
@@
This expression is the generalized Euler equation that we were seeking. The additional correction involves a derivative of the pressure $p$ with respect to $y$. Therefore, if the thermodynamics of a fluid in a given spacetime does not respond to variations in $y$, the additional correction $\frac{\partial p}{\partial y}$ vanishes recovering the standard Euler identity.
Another important aspect of the generalized effective action for a perfect fluid (40) is the emergence of a new conserved current. In addition to the off-shell conservation of the entropy current (37) and the on-shell conservation of the charge current (38), we now also have:
$\nabla_{a}J_{(y)}^{a}=0,\quad J_{(y)}^{a}=u^{a}\partial_{y}p~{}.$ (48)
This conservation law arises from the fact that the variable $y$, defined in (39), is invariant under spatial diffeomorphisms of the target space, $\sigma^{I}\to\sigma^{I}+f^{I}(\sigma^{J})$. Mathematically, the pair $(y,\partial_{y}p)$ is analogous to the chemical potential $\mu$ and the charge density $q$ associated with the $U(1)$ symmetry. However, it is important to emphasize that the variable $y$ was introduced to capture a geometric scale imposed by the underlying spacetime on the thermodynamics of the fluid. For this reason, unlike conventional thermodynamic variables, $y$ is not expected to admit a microscopic interpretation.
To further explore the meaning of the variable $y$, let us consider a scenario of global thermal equilibrium, where the fluid clock can be aligned with physical time; that is, $\sigma^{0}=x^{0}$. In this case, we find that $y=u^{0}$, a result that may appear somewhat opaque at first glance. To build intuition, let us perform a Weyl transformation of the physical metric, $g^{\prime}=Wg$, where $W$ is a conformal factor. Under this transformation, the fluid velocity transforms as $u^{\prime a}=\frac{u^{a}}{W}$. This leads to the following relation:
$y=u^{\prime 0}=\frac{u^{0}}{W}~{}.$ (49)
This expression suggests that the variable $y$ is naturally associated with the inverse of the conformal factor $W$. Alternatively, rather than performing an explicit Weyl transformation, one can express the physical metric as a conformal rescaling of a reference metric, $g=W\bar{g}$, leading to the same conclusion that $y\sim\frac{1}{W}$.
To facilitate comparison between the central result of this work and existing literature, we rewrite the generalized Euler identity (47) in the following form:
@@
where $\omega$ denotes the density of the grand canonical potential. This expression highlights that the presence of an additional scale in the system, encoded by the variable $y$, challenges the conventional assumption that, in the thermodynamic limit, the grand canonical potential $\omega$ is simply proportional to the pressure $p$.
The relation in (50) bears a notable resemblance to earlier results in the literature on effective actions for fluids Minwalla ; Jensen0  [^4] We thank Kristan Jensen for bringing these references to our attention.. However, a key distinction lies in the construction of our effective action, which is formulated at zeroth order in the derivative expansion. In contrast, the works cited in Minwalla ; Jensen0 incorporate higher-order derivative corrections, which modify the form of the free energy accordingly.
@@ 4 Smarr formula as a particular case of the generalized Euler equation
In this section, we extend the discussion from section 2 by examining the AdSd Reissner-Nordström black hole with arbitrary horizon topology. We adopt Eddington-Finkelstein coordinates, as this gauge simplifies the analysis of Weyl transformations compared to the Fefferman-Graham gauge. Moreover, we restrict the boundary dimension $(d-1)$ to be odd; otherwise, the conformal anomaly and the Casimir energy would need to be accounted for in the effective action. The line element is:
@@
$M+pV=TS+\mu Q+\frac{k\Sigma_{k,d-2}}{8\pi G_{N}}r^{d-3}_{h}~{}.$ (55)
For for $k=0$ the last term vanishes, thereby recovering the standard Euler equation for a charged black brane. This is because, in flat space, there are no additional scales to consider.
We will now compute the thermodynamic potential $\chi$ (40) which is associated with the effective action for the holographic perfect fluid dual to the AdSd Reissner-Nordström black hole. Our objective is to place the holographic fluid in a boundary spacetime that exhibits spacetime dependence, allowing us to mimic a state that is near out of equilibrium [^5] This step is purely for illustrative purposes and carries no additional meaning, as we are disregarding dissipation, and anomalies.. As a consequence, the thermodynamic variables now will become local functions of the spacetime. To obtain the boundary metric with the desired properties, we perform the following radial diffeomorphism in the bulk line element (51):
@@
This line element describes a non-homogeneous, time-dependent background in which the fluid resides. The radial diffeomorphism (56) consequently induces a Weyl transformation on the boundary metric [^6] In other words, the Weyl transformation is induced by a PBH transformation of the bulk metric. A detailed discussion of the consequences of this transformation in the thermodynamics of asymptotically locally AdS black holes is given in Ioannis3 ., leading to the following expression for the fluid velocity:
$u_{a}=-\frac{W}{l}\delta^{0}_{a}~{}.$ (60)
@@
$ds^{2}_{fluid}=-0\cdot du^{2}+r^{2}_{h}d\Sigma^{2}_{k}~{},$ (61)
that matches the horizon metric  [^7] More precisely, the horizon metric before the Weyl transformation.. Given the target metric and the fluid velocity, we can use the formulas (29), (31), and (39) to compute the entropy density $s$, the chemical potential $\mu$, and the variable $y$, respectively:
$s=\frac{\Sigma_{k,d-2}}{4G_{N}}\frac{r_{h}^{d-2}}{W^{d-2}}~{},\quad\mu=\frac{\tilde{q}}{\gamma r_{h}^{d-3}}\frac{1}{W}~{},\quad y=\frac{l}{W}~{},$ (62)
which are the variables that the thermodynamic potential $\chi$ depends on. We note that the entropy density satisfies its conservation law (37) as expected. We will utilize these variables to derive the explicit expression for the new thermodynamic potential $\chi$ from the energy density $\epsilon$ using equation (42). Our next step involves computing the boundary stress tensor via the formula (12) and the boundary metric given in (59). The components of the stress tensor are
$-T^{t}_{\;t}=\frac{lM}{W^{d-1}\Sigma_{k,d-2}}~{},$ (63)
$T^{\phi_{j}}_{\;\phi_{j}}=\frac{lM}{(d-2)W^{d-1}\Sigma_{k,d-2}}~{}.$ (64)
@@
$\epsilon=\frac{(d-2)}{16\pi G_{N}l}\left[\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l)^{2}\mu^{2})+\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-1}{d-2}}\right]~{},$ (65)
where $M$ has been expressed in terms of $r_{h}$ as an intermediate step, and then in terms of the variables $s$, $\mu$, and $y$ given in equation (62). One can also verify that the conformal equation of state: $p=\frac{\epsilon}{d-2}$ is satisfied.
@@
$\chi=\frac{(d-2)}{16\pi G_{N}l}\left[-\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l)^{2}\mu^{2})+\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-1}{d-2}}\right]~{}.$ (67)
@@
$T=\frac{(d-1)}{4\pi l}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{1}{d-2}}+\frac{(d-3)(ky^{2}-(\gamma l)^{2}\mu^{2})}{4\pi l}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{-1}{d-2}}~{},$ (68)
$q=\frac{(d-2)\gamma^{2}l\mu}{8\pi G_{N}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}~{}.$ (69)
@@
$\frac{\partial p}{\partial y}=\frac{ky}{8\pi G_{N}l}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}~{}.$ (70)
Consequently, this term vanishes for a charged black brane, as expected, since flat spacetime does not introduce any additional geometrical scale. Through straightforward algebraic manipulations, one can verify that the generalized Euler equation holds:
@@
Furthermore, substituting the expressions for $y$ and $\mu$ from (62), one can confirm that $q$ [^8] Let us emphasize that $\tilde{q}$ denotes the constant appearing in the black hole line element, whereas $q=q(x)$ is a local function defined on the boundary spacetime. We apologize for any notational inconvenience this may cause. and $\partial_{y}p$ satisfy their respective conservation laws, given by (38) and (48).
@@
In conclusion, we have demonstrated that the Smarr formula is a specific instance of the generalized Euler equation (71). This result implies that holographic fluids satisfying the Smarr formula are not fundamentally distinct from more conventional fluids. The key insight is that when an additional scale arises from the underlying geometry of the thermal system, the standard scaling argument used to derive the Euler equation no longer holds, necessitating a generalized formulation. For a system with a compact volume, the physical interpretation is quite clear: the correction term $y\frac{\partial p}{\partial y}$ in the Euler equation naturally accounts for a finite-volume effect, as previously noted by Karch . In summary, there is nothing intrinsically special about the Smarr formula for AdS black holes; it emerges naturally from general thermodynamic and fluid considerations.
5 Comments on cosmological constant as a thermodynamic variable
In the previous sections, we argued and demonstrated that the Smarr formula is a specific case of the generalized Euler equation, particularly in the context of black holes in AdS. In this sense, the Smarr formula naturally fits within the framework of the AdS/CFT correspondence, as its origins are rooted in well-established physics. This conclusion is supported by the fact that we derived the generalized Euler equation using effective field theory techniques applied to fluids. Only after establishing this result did we invoke the AdS/CFT correspondence to interpret the Smarr formula.
@@ 5 Comments on cosmological constant as a thermodynamic variable
where $d$ is the dimension of the bulk spacetime [^9] For reviews on extended black hole thermodynamics, see Mann4 ; Mann5 . Other relevant literature on the subject includes Mann2 ; Mann3 ; Clifford2 ; Kastor1 .. Clearly, the notion of pressure $p_{\Lambda}$ does not correspond to the pressure of the dual field theory. For this reason, the physical interpretation of the Smarr formula was investigated from a holographic perspective in Karch .
@@
The main result of this paper—that the Smarr formula is a specific instance of the generalized Euler equation—conflicts this new interpretation for extended black hole thermodynamics. The reasoning is as follows: perfect fluids, a particular case of hydrodynamics, differs fundamentally from renormalization group (RG) physics. While the number of degrees of freedom $N$ can indeed change with the energy scale in RG flow (as is well understood in the context of AdS/CFT, see Ioannis1 ; Ioannis2 ), this cannot explain the Smarr formula. In a perfect fluids, thermodynamic variables must be functions of spacetime to describe local thermal equilibrium, something that the number of degrees of freedom $N$ cannot do. The point is that we should not confuse a flow in a physical space where the fluid is on with a flow in theory space where an RG flow takes place. Finally, it is crucial to recognize that global equilibrium is merely a special case of local equilibrium. As a result, there is no justification for treating $C\sim N^{2}$ as a thermodynamic variable.
@@ 6 Discussion
In this work, we derived a generalized Euler equation (47) using effective field theory techniques, granting our result a broader range of applicability independent of the AdS/CFT correspondence. We then applied the generalized Euler equation to a specific case: a holographic fluid derived from the boundary stress tensor associated with an AdSd Reissner-Nordström black hole, from which we recovered the Smarr formula. Based on these findings, we further argue that treating the cosmological constant as a thermodynamic variable is not physically justified. This is because any thermodynamic variable associated with the AdS radius $l$ cannot become a function of the boundary spacetime, as required for local thermodynamic equilibrium.
@@
It has been demonstrated that evaluating the on-shell action of spacetimes with a future event horizon (featuring a planar horizon topology) can reproduce the standard effective action presented in (32). This was achieved by setting up a double-Dirichlet boundary problem to evaluate the on-shell action Natalia1 ; Liu1 . It would be intriguing to extend this construction to the more general effective action (40) discussed in our article. For this purpose, it is necessary to gain a deeper understanding of the variable $y$ introduced in this work.
@@
Another intriguing avenue for exploration is the explicit construction of the thermodynamic potential $\chi$ for rotating black holes, which also exhibit a boundary stress tensor with a perfect fluid form Clifford0 ; Clifford . In this work, we focused on even boundary dimensions $d$ to avoid the complications arising from conformal anomalies. It would be valuable to develop the most general effective action for perfect fluids that can incorporate such conformal anomalies. In lower-dimensions, a notable example is the Schwarzian action, which can be interpreted as the effective action of an anomalous perfect fluid in 1d Jensen2 . A similar structure appears in the boundary action of a non-rotating BTZ black hole with external sources Jensen3 ; Kevin ; Me .
It’s well-known that the Schwarzian action captures the near-horizon dynamics of near-extremal black holes Amhed ; Stanford . In contrast, our effective action given in (40) is an off-shell formulation that captures not only the one-point functions (the stress tensor), but also the Ward identities (the Navier-Stokes equations without dissipation). In this way, following closely the work of Me , it should be possible to derive a decoupling limit from the action (67), which encodes the dynamics of a holographic charged fluid, and connect it to the one-dimensional anomalous perfect fluid of Jensen2 . In fact, the author of Me already achieved this decoupling limit for a restricted class of CFT3, although using a different set of variables from the fluid variables.
The Smarr formula remains valid in gravity theories with higher-order curvature corrections Tomas1 . This raises an important question: can the generalized Euler equation (47) accommodate these theories? If not, it would suggest that the Smarr formula once again indicates the need for further generalizations of the Euler equation. Similar questions arise in the context of the dyonic AdS black holes with a non-trivial scalar profile discussed in Sugra1 ; Sugra2 , which constitute solutions to gauged supergravity.
@@
It is worth emphasizing that the isometries of the target metric $G_{MN}$, as defined in (24), correspond to the emergent low-energy gauge symmetries of the hydrodynamic effective action. More broadly, asymptotic symmetries at the horizon ASG1 ; ASG2 ; ASG3 ; ASG4 ; ASG5 can be interpreted as manifestations of these emergent gauge symmetries of the boundary theory. Consequently, any generalization of horizon symmetries implies the presence of new gauge symmetries in holographic effective actions. Exploring potential connections between the present work and the novel symmetries identified in Liu5 would be a compelling direction for future research.